A truck on a straight road starts from rest, accelerating at 2.00 m/s2 until it reaches a speed of 20.0 m/s. Then the truck travels for 20 s at a constant speed until the brakes are applied, stopping the truck in a uniform manner in an additional 5.00 s. (a) How long is the truck in motion? (b) What is the average velocity of the truck for the motion described?
1
Expert's answer
2020-10-08T18:13:49-0400

Let's determine the travel time and distance for each section of the path

1) For the first leg of the path

"t_1=\frac{v}{a_1}=\frac{20}{2}=10s"

"s_1=\frac{a\cdot {t_1}^2}{2}=\frac{2\cdot 10^2}{2}=100m"

2)For the second leg

"t_2=20s"

"s_2=v \cdot t_2=20 \cdot 20=400m"

3)For the third leg

"t_3=5s"

Determine the acceleration with which the truck slows down

"a_3=\frac{v}{t_3}=\frac{20}{5}=4m/s^2"

then

"s_3=v \cdot t_3-\frac{a_3\cdot {t_3}^2}{2}=20 \cdot 5-\frac{4\cdot 5^2}{2}=50m"

a)the truck is in motion

"t=t_1+t_2+t_3=10+20+5=35s"

b)average speed of the truck during the described movement

"v_m= \frac{s_1+s_2+s_3}{t_1+t_2+t_3}=\frac{100+400+50}{10+20+5}=15.714m/s"